# Stochastic SIR model with individual heterogeneity and infection-age dependent infectivity on large non-homogeneous random graphs

### v1

@@ 2.1. Infection-age dependent infectivity
where $F^{c}_{X^{N}_{j}}(A^{N}_{j}(0))>0$ holds a.s. for any $j\in\mathcal{I}^{N}(0)$. Moreover, the function $\chi:\mathbf{\mathbb{X}}\times\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ defined as $\chi(x,a)=\bar{\lambda}(x,a)/F^{c}_{x}(a)$ if $F^{c}_{x}(a)>0$ and $\chi(x,a)=0$ otherwise, is upper-bounded by $\lambda_{*}$.
Recall that $\mathbb{E}_{0}^{N}\big{[}\lambda^{N}_{i}(a)\big{]}$ equals by definition $\mathbb{E}\big{[}\lambda^{N}_{i}(a)\,\big{|}\,\mathcal{F}^{N}_{0}\big{]}$ and similarly for $\mathbb{E}_{0}^{N}\big{[}\lambda^{N,0}_{j}(t)\big{]}$.
@@
[…]
and is actually upper-bounded by $\lambda_{*}$ as soon as $\widehat{\lambda}$ is itself upper-bounded by $\lambda_{*}$.
@@
In this work we generally allow for intricate dependencies between the possible value of $\lambda_{i}^{N}(a)$ and the infection duration of individual $i$. Though, a classical choice for the function $\widehat{\lambda}$ is to be defined in terms of two given deterministic functions $\widetilde{\lambda}(x,a)$ and $g(x,y)$ as follows: $\widehat{\lambda}(x,y,a)=\widetilde{\lambda}(x,a)\mathbbm{1}_{\left\{a<g(x,y)\right\}}$. In this case, the expression for $\bar{\lambda}$ simplifies as follows:
$\bar{\lambda}(X_{i}^{N},a)=\widetilde{\lambda}(X_{i}^{N},a)\cdot F^{c}_{X^{N}_{i}}(a).$
@@ 2.2. Construction of the connectivity graph
The typical construction of such a random graph starts from a graphon kernel. For example, assuming here that $\mathbf{\mathbb{X}}=[0,1]$, let $\bar{\kappa}(x,x^{\prime})$ be a deterministic function on $[0,1]^{2}$, such as $\bar{\kappa}(x,x^{\prime})=x\cdot x^{\prime}$. One can sample a $N$-node random graph with $\kappa^{N}=\bar{\kappa}$ so that $\mathbb{P}_{0}^{N}\big{(}i\stackrel{{\scriptstyle N}}{{\sim}}j\big{)}=\bar{\kappa}(X^{N}_{i},X^{N}_{j})$ for nodes $i,j=1,\dots,N$, where the $X^{N}_{i}$ are uniformly and independently sampled in $[0,1]$. Another very natural choice of $X^{N}_{i}$ is also $X^{N}_{i}=i/N$ for $i=1,\dots,N$, which also leads to $\bar{\mu}_{X}$ being the Lebesgue measure on $[0,1]$. In the case $\bar{\kappa}(x,x^{\prime})=x\cdot x^{\prime}$, the degree of individual $i$ in the graph then scales linearly with $N$ and with the individual type $X_{i}^{N}$. The fact that the degree scales linearly in $N$ is typically how a “dense” graph is defined. Yet, we allow for more generality in the graph density with a sequence $(\kappa^{N})$ of functions that may scale with $N$. The first way to do so is to introduce a scaling factor $\varepsilon^{N}>0$, typically $\varepsilon^{N}=N^{-\alpha}$ with $\alpha\in(0,1)$, so that $\kappa^{N}=\varepsilon^{N}\cdot\bar{\kappa}$ as in [37]. Following [20], this form of scaling is however not assumed to allow for even more general dependencies between the pair of individual characteristics $(X_{i}^{N},X_{j}^{N})$ and $N$. The role of the denseness of the graph will be discussed in Subsection 3.1, after the statement of our main theorem.
@@ 2.3. Epidemic dynamics
Note that heterogeneity in the susceptibility to infection can be included in the model through this values $\omega^{N}(i,j)$, such heterogeneity between individuals being possibly partly explained by the characteristic $X_{i}^{N}$. This is a major reason for us not to assume that $\omega^{N}$ is symmetric. Such a framework for $\omega^{N}(i,j)$ allows in addition more realistic and intricate relations between the contact rates of individuals $i$ and $j$ depending on their respective characteristics $X_{i}^{N}$ and $X_{j}^{N}$. Even more generally than contact matrices typically exploited for epidemiological predictions, see, e.g., [43], we allow for the interplay between characteristics that are continuous and for an additional degree of independent randomness.
The values $(\omega^{N}(i,j))_{i,j\in[\![1,N]\!]}$ are assumed in addition to be mutually independent between different edges and independent of the $(\lambda_{i}^{N})_{i\leq N}$ conditionally on $\mathcal{F}_{0}^{N}$, that is, of the $(Y_{i}^{N})$ if we consider the formalism in Remark 2.1. The deterministic measurable function $\gamma^{N}:\mathbf{\mathbb{X}}\times\mathbf{\mathbb{X}}\to\mathbb{R}_{+}$ captures the expected infectious rate of interaction of active contacts, in that the following identity is assumed:
@@
We may also allow for a degree of variability in the infectious rate of interaction, that we synthetize with the following variance term $\upsilon^{N}$, for any $(i,j)\in\mathcal{E}^{N}$:
Given the above definition of $\gamma^{N}$, this definition of $\upsilon^{N}$ shall be interpreted as follows:
Aa a consequence of these definitions, we have
@@
so that the infection time $\tau_{i}^{N}$ is the unique jump time of $D_{i}^{N}$ (which takes the value $\infty$ in the absence of such a jump), while $\mathcal{D}^{N}(t)$ is the subset of initially susceptible individuals infected by the disease by time $t$ after time 00, or equivalently those $i\in\mathcal{S}^{N}(0)$ such that $D_{i}^{N}(t)=1$. Note that $A_{i}^{N}(t)>0$ means $D_{i}^{N}(t^{-})=1$, which is not exactly $D_{i}^{N}(t)=1$. In other words, $\mathcal{D}^{N}(t)=(\mathcal{I}^{N}(t)\cup\mathcal{R}^{N}(t))\setminus(\mathcal{I}^{N}(0)\cup\mathcal{R}^{N}(0))$.
@@ 3. Functional law of large numbers
There exist finite measures $\bar{\mu}^{S}_{0}(\mathrm{d}x)$ on $\mathbf{\mathbb{X}}$, $\bar{\mu}^{I}_{0}(\mathrm{d}x,\mathrm{d}a)$ on $\mathbf{\mathbb{X}}\times{\mathbb{R}}_{+}$ and $\bar{\mu}^{R}_{0}(\mathrm{d}x)$ on $\mathbf{\mathbb{X}}$ that are the weak limits in probability as $N\to\infty$ of respectively $\bar{\mu}_{0}^{S,N}$, $\bar{\mu}_{0}^{I,N}$, and $\bar{\mu}_{0}^{R,N}$. In addition, the following inequality holds between the measures $\bar{\mu}^{R}_{0}$ and $\bar{\mu}^{I}_{0}$:
$\bar{\mu}^{R}_{0}(\mathrm{d}x)\geq\int_{0}^{\infty}\frac{F_{x}(a)}{F_{x}^{c}(a)}\bar{\mu}^{I}_{0}(\mathrm{d}x,da).$ (3.2)
@@
Remark 3.2.
The property that relates $\bar{\mu}^{R}_{0}$ to $\bar{\mu}^{I}_{0}$ in Assumption 3.1 is justified by the following guiding idea: individuals infected with age $a$ at time 00 should be interpreted as the outcomes of infection events at time $-a$ that produced a larger proportion of infected (with a scaling factor $[F_{x}^{c}(a)]^{-1}$), from which a proportion $F_{x}(a)$ got to recover.
In addition, we make the following two assumptions to capture the behavior of the graph structure $(\omega^{N}(i,j))_{i,j\in[\![1,N]\!]}$ as $N$ tends to infinity, firstly in terms of conditional expectations.
@@
Complementary to Assumptions 2.4 and 2.6, the function $:(x,a)\mapsto F_{x}(a)$ is $\bar{\mu}^{I}_{0}$-a.e. continuous while the function $:x\mapsto\bar{\lambda}(x,.)$ taking values in $\mathcal{D}(\mathbb{R}_{+},\mathbb{R}_{+})$ equipped with the uniform norm topology is $\bar{\mu}_{X}$-a.e. continuous.
The sequence $\bar{\omega}^{N}$ is uniformly upper-bounded by virtue of Assumption 3.3:
Lemma 3.4.
Under Assumption 3.3, there exists a constant $\omega^{*}>0$ that is an upper-bound on $\bar{\omega}^{N}$ uniformly on $\mathbf{\mathbb{X}}^{2}$ and for $N$ sufficiently large.
Without loss of generality regarding our convergence results in $N\to\infty$, we assume in the following that $\omega^{*}>0$ is actually an upper-bound uniformly on $\mathbf{\mathbb{X}}^{2}$ and on any $N$.
@@
To be clear with our notations of conditional expectations, $\mathbb{E}_{0}^{N}\big{[}\upsilon^{N}(i,j)\,\Big{|}\,(i,j)\in\mathcal{E}^{N}\big{]}$ describes the ratio of $\mathbb{E}\big{[}\upsilon^{N}(i,j);(i,j)\in\mathcal{E}^{N}\,\big{|}\,\mathcal{F}_{0}^{N}\big{]}$ by $\mathbb{P}\big{[}(i,j)\in\mathcal{E}^{N}\,\big{|}\,\mathcal{F}_{0}^{N}\big{]}=\kappa^{N}(X_{i}^{N},X_{j}^{N})$.
For brevity, we denote $\mathcal{D}_{1}:=\mathcal{D}({\mathbb{R}}_{+},\mathcal{M}(\mathbf{\mathbb{X}}))$ and $\mathcal{D}_{2}:=\mathcal{D}({\mathbb{R}}_{+},\mathcal{M}(\mathbf{\mathbb{X}}\times\mathbb{R}_{+}))$.
@@
$\begin{split}\overline{\mathfrak{F}}(t,x)&=\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\bar{\mu}_{0}^{I}(\mathrm{d}x^{\prime},da^{\prime})\\ &\quad+\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\bar{\lambda}(x^{\prime},t-s)\overline{\mathfrak{F}}(s,x^{\prime})\bar{\mu}^{S}_{s}(\mathrm{d}x^{\prime})\mathrm{d}s\,,\end{split}$ (3.7)
with $\bar{\lambda}$, $\bar{\omega}$ being given respectively in Assumptions 2.6 and 3.3. The uniqueness of the above system is stated among the potential candidates $(\bar{\mu}^{S}_{t},\overline{\mathfrak{F}}(t))_{t\geq 0}$ in which $\bar{\mu}^{S}_{\cdot}$ is a $\mathcal{M}(\mathbf{\mathbb{X}})$-valued càd-làg process such that $\mu_{t}(\mathbf{\mathbb{X}})\leq 1$ for any $t\geq 0$, while $\overline{\mathfrak{F}}(\cdot)$ is a càd-làg process whose values are bounded measurable functions from $\mathbf{\mathbb{X}}$ to $\mathbb{R}_{+}$, with local in time upper-bounds.
Given the pair $(\bar{\mu}^{S}_{\cdot},\overline{\mathfrak{F}}(\cdot))$, $\langle\bar{\mu}^{I}_{t},\psi\rangle$ and $\langle\bar{\mu}^{R}_{t},\varphi\rangle$ are given as follows, with the test functions $\psi\in C_{b}(\mathbf{\mathbb{X}}\times{\mathbb{R}}_{+})$ and $\varphi\in C_{b}({\mathbb{R}}_{+})$,
$\begin{split}\langle\bar{\mu}^{I}_{t},\psi\rangle&=\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\psi(x,a+t)\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)\\ &\quad+\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\psi(x,t-s)F^{c}_{x}(t-s)\overline{\mathfrak{F}}(s,x)\bar{\mu}^{S}_{s}(\mathrm{d}x)\mathrm{d}s\,,\end{split}$ (3.8)
@@
$\begin{split}\langle\bar{\mu}^{R}_{t},\varphi\rangle&=\langle\bar{\mu}^{R}_{0},\varphi\rangle+\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\varphi(x)\bigg{(}1-\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bigg{)}\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)\\ &\quad+\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\varphi(x)F_{x}(t-s)\overline{\mathfrak{F}}(s,x)\bar{\mu}^{S}_{s}(\mathrm{d}x)\mathrm{d}s\,.\end{split}$ (3.9)
Note that the unique solution to (3.6) is given by
$\bar{\mu}^{S}_{t}(\mathrm{d}x)=\,\exp\bigg{(}-\int_{0}^{t}\overline{\mathfrak{F}}(s,x)\mathrm{d}s\bigg{)}\bar{\mu}^{S}_{0}(\mathrm{d}x)\,.$ (3.10)
3.1. Discussions of the main result
@@
Our conditions on the limiting kernel of interaction $\bar{\omega}$ encompasses most, if not all, of the pairwise kernel interactions. Contact matrices over a discrete space are allowed, as, e.g., in [11], as well as spatial interactions as considered in [36] with characteristics that are distributed on a compact subset of ${\mathbb{R}}^{d}$, or even combination of both settings. A possible extension of our result would be to consider individual types that evolves in time, typically when the individual locations move in space. It is another natural objective to establish the FLLN in instances where the limiting kernel $\bar{\omega}$ incorporates an additional dependency on the overall distribution $\bar{\mu}_{X}$ as in the general setting of [36].
@@
Let us recall the previously introduced example where $\bar{\kappa}^{N}=\bar{\kappa}$ for any $N$ and $\bar{\kappa}(x,x^{\prime})=x\cdot x^{\prime}$ while the $(X_{i}^{N})$ are uniformly sampled in $[0,1]$. Let us assume in addition that the contact rate function $\gamma^{N}$ is fixed and does not depend on the types of the individuals $i$ and $j$ in contact while $\bar{\omega}$ is non-zero. In this case, $\Upsilon^{N}\equiv 0$.
@@
$\check{\mathfrak{F}}(t)=\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}x^{\prime}\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\bar{\mu}_{0}^{I}(\mathrm{d}x^{\prime},da^{\prime})+\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}x^{\prime}\cdot\bar{\lambda}(x^{\prime},t-s)\overline{\mathfrak{F}}(s,x^{\prime})\bar{\mu}^{S}_{s}(\mathrm{d}x^{\prime})\mathrm{d}s\,.$
As considered in [37], such a setting for the dynamics on a graphon can be generalized by assuming a scaling factor $\varepsilon^{N}>0$ depending on $N$, for instance $\varepsilon^{N}=N^{-\alpha}$ with $\alpha\in(0,1)$, in that $\bar{\kappa}^{N}=\varepsilon^{N}\cdot\bar{\kappa}$. Provided we again assume the contact rate to be fixed constant at a value $\bar{\gamma}^{N}>0$, we can follow the same reasoning except that $\bar{\gamma}^{N}$ is then equivalent to $\bar{\gamma}/(N\varepsilon^{N})$ instead of $\bar{\gamma}/N$ and that $N\varepsilon^{N}\to\infty$ is required (and sufficient) for Assumption 3.5 to hold.
Note also that $N\varepsilon^{N}\to\infty$ typically means that the node degrees are asked to go to infinity, without condition on the relation to the number $N$ of nodes. We refer to [20] for more details on the relation to the denseness level of the graph (in terms notably of the number of edges or the node degrees). The additional condition proposed in [37] corresponding to the convergence to infinity of $N\varepsilon^{N}/\log(N)$ appears related to their technique of proof rather than to the convergence of the epidemic process itself. The case where $\varepsilon^{N}=N^{-1}$ leads to limiting equations of a different nature, as hinted with exploratory simulations in [20] and explicited in the result of [10] for instances of stochastic block models.

### v2

@@ 2.1. Infection-age dependent infectivity
where $F^{c}_{X^{N}_{j}}(A^{N}_{j}(0))>0$ holds a.s. for any $j\in\mathcal{I}^{N}(0)$. Moreover, the function $\chi:\mathbf{\mathbb{X}}\times\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ defined as $\chi(x,a)=\bar{\lambda}(x,a)/F^{c}_{x}(a)$ if $F^{c}_{x}(a)>0$ and $\chi(x,a)=0$ otherwise, is upper-bounded by $\lambda^{*}$.
Additional conditions on the regularity of $\bar{\lambda}$ will be stated in Assumption 3.2, once the limiting measures on $\mathbf{\mathbb{X}}$ will be defined. Recall that $\mathbb{E}_{0}^{N}\big[\lambda^{N}_{i}(a)\big]$ equals by definition $\mathbb{E}\big[\lambda^{N}_{i}(a)\,\big|\,\mathcal{F}^{N}_{0}\big]$ and similarly for $\mathbb{E}_{0}^{N}\big[\lambda^{N,0}_{j}(t)\big]$.
@@
$\frac{\bar{\lambda}(x,a)}{F^{c}_{x}(a)}=\frac{\int_{\mathbf{\mathbb{Y}}}\widehat{\lambda}(x,y,a)\,\mathbbm{1}_{\left\{g(x,y)>a\right\}}\;\mu_{Y}(\mathrm{d}y)}{\int_{\mathbf{\mathbb{Y}}}\mathbbm{1}_{\left\{g(x,y)>a\right\}}\;\mu_{Y}(\mathrm{d}y)}\,$,
and is actually upper-bounded by $\lambda^{*}$ as soon as $\widehat{\lambda}$ is itself upper-bounded by $\lambda^{*}$.
@@
In this work, we generally allow for intricate dependencies between the possible value of $\lambda_{i}^{N}(a)$ and the infection duration of individual $i$. Though, a classical choice for the function $\widehat{\lambda}$ is to be defined in terms of two given deterministic positive functions $\widetilde{\lambda}(x,a)$ and $g(x,y)$ as follows: $\widehat{\lambda}(x,y,a)=\widetilde{\lambda}(x,a)\mathbbm{1}_{\left\{a<g(x,y)\right\}}$. In this case, the expression for $\bar{\lambda}$ simplifies as follows:
$\bar{\lambda}(X_{i}^{N},a)=\widetilde{\lambda}(X_{i}^{N},a)\cdot F^{c}_{X^{N}_{i}}(a).$ (2.3)
@@
Remark 2.10.
In the previous assumptions, we did not allow the infectious period to extend indefinitely, in line with the SIR formalism. It does however not appear crucial in the following proofs. Some SI formalism without recovery or even a mix between SI and SIR formalism could be treated very similarly. We mean situations where $\mathbb{P}(\eta^{N}_{i}=+\infty)>0$ for some $i$, with possibly $F_{x}$ upper-bounded by a constant strictly smaller than 1 for some non-negligible set of $x\in\mathbf{\mathbb{X}}$ and the function $g$ in Remark 2.5 with possibly infinite values.
@@ 2.2. Construction of the connectivity graph
Remark 2.11.
The typical construction of such a random graph starts from a graphon kernel. For example, assuming here that $\mathbf{\mathbb{X}}=[0,1]$, let $\bar{\kappa}(x,x^{\prime})$ be a deterministic function on $0,1]^{2}$, such as $\bar{\kappa}(x,x^{\prime})=x\cdot x^{\prime}$. One can sample a $N$-node random graph with $\kappa^{N}=\bar{\kappa}$ so that $\mathbb{P}_{0}^{N}\big(i\stackrel{{\scriptstyle N}}{{\sim}}j\big)=\bar{\kappa}(X^{N}_{i},X^{N}_{j})$ for nodes $i,j=1,\dots,N$, where the $X^{N}_{i}$ are uniformly and independently distributed in $[0,1]$. Another very natural choice of $X^{N}_{i}$ is also $X^{N}_{i}=i/N$ for $i=1,\dots,N$, which also leads to the limiting empirical measure $\bar{\mu}_{X}$, as defined in the formula (3.3) below, being the Lebesgue measure on $[0,1]$.
Remark 2.12.
In the case $\bar{\kappa}(x,x^{\prime})=x\cdot x^{\prime}$, the degree of individual $i$ in the graph then scales linearly with $N$ and with the individual type $X_{i}^{N}$. The fact that the degree scales linearly in $N$ is typically how a “dense” graph is defined. Yet, we allow for more generality in the graph density with a sequence $(\kappa^{N})$ of functions that may scale with $N$. The first way to do so is to introduce a scaling factor $\varepsilon^{N}>0$, typically $\varepsilon^{N}=N^{-\alpha}$ with $\alpha\in(0,1)$, so that $\kappa^{N}=\varepsilon^{N}\cdot\bar{\kappa}$ as in [35]. Following [18], this form of scaling is however not assumed in order to allow for even more general dependencies between the pair of individual characteristics $(X_{i}^{N},X_{j}^{N})$ and $N$. The role of the denseness of the graph will be discussed in Section 3, after the statement of our main theorem.
@@ 2.3. Epidemic dynamics
Remark 2.13.
Heterogeneity in the susceptibility to infection can be included in the model through this values $\omega^{N}(i,j)$, such heterogeneity between individuals being possibly partly explained by the characteristic $X_{i}^{N}$. This is a major reason for us not to assume that $\omega^{N}$ is symmetric. Such a framework for $\omega^{N}(i,j)$ allows in addition more realistic and intricate relations between the contact rates of individuals $i$ and $j$ depending on their respective characteristics $X_{i}^{N}$ and $X_{j}^{N}$. Even more generally than contact matrices typically exploited for epidemiological predictions, see, e.g., [40], we allow for the interplay between characteristics that are continuous and for an additional degree of independent randomness.
Assumption 2.14.
The values $(\omega^{N}(i,j))_{i,j\in[\![1,N]\!]}$ are mutually independent between different edges and independent of the $(\lambda_{i}^{N})_{i\leq N}$ conditionally on $\mathcal{F}_{0}^{N}$. There exists a deterministic measurable function $\gamma^{N}:\mathbf{\mathbb{X}}\times\mathbf{\mathbb{X}}\to\mathbb{R}_{+}$ such that:
@@
Remark 2.15.
Under the formalism given in Remark 2.1, the values $(\omega^{N}(i,j))_{i,j\in[\![1,N]\!]}$ are independent of the $(\lambda_{i}^{N})_{i\leq N}$ if they are independent of the $(Y_{i}^{N})$, conditionally on $\mathcal{F}_{0}^{N}$.
The function $\gamma^{N}$ captures the expected infectious rate of interaction for active contacts. As a consequence of these definitions, we have
@@
We also allow for a degree of variability in the infectious rate of interaction, that we synthetize with the following variance term $\upsilon^{N}$, for any $(i,j)\in\mathcal{E}^{N}$:
Our upcoming Assumption 3.4 entails that the quantity $\upsilon^{N}(i,j)$ is finite, for any $i,j\in[\![1,N]\!]$. Given the above definition of $\gamma^{N}$, this definition of $\upsilon^{N}$ shall be interpreted as follows:
@@
so that the infection time $\tau_{i}^{N}$ is the unique jump time of $D_{i}^{N}$ (which takes the value $\infty$ in the absence of such a jump), while $\mathcal{D}^{N}(t)$ is the subset of initially susceptible individuals infected by the disease by time $t$ after time 0, or equivalently those $i\in\mathcal{S}^{N}(0)$ such that $D_{i}^{N}(t)=1$. In other words, $\mathcal{D}^{N}(t)=(\mathcal{I}^{N}(t)\cup\mathcal{R}^{N}(t))\setminus(\mathcal{I}^{N}(0)\cup\mathcal{R}^{N}(0))$.
@@ 3. Functional law of large numbers
There exist finite measures $\bar{\mu}^{S}_{0}(\mathrm{d}x)$ on $\mathbf{\mathbb{X}}$, $\bar{\mu}^{I}_{0}(\mathrm{d}x,\mathrm{d}a)$ on $\mathbf{\mathbb{X}}\times{\mathbb{R}}_{+}$ and $\bar{\mu}^{R}_{0}(\mathrm{d}x)$ on $\mathbf{\mathbb{X}}$ that are the weak limits in probability as $N\to\infty$ of respectively $\bar{\mu}_{0}^{S,N}$, $\bar{\mu}_{0}^{I,N}$, and $\bar{\mu}_{0}^{R,N}$.
@@
In addition, we make the following two assumptions to capture the behavior of the graph structure $(\omega^{N}(i,j))_{i,j\in[\![1,N]\!]}$ as $N$ tends to infinity, firstly in terms of conditional expectations. We include additional regularity conditions together with the first assumption, on the quantities introduced in Assumptions 2.4, 2.6 and 3.1.
@@
$\bar{\mu}^{I}_{0}$ is absolutely continuous with respect to $(\bar{\mu}_{X}\otimes Leb)$, where $Leb$ denotes the Lebesgue measure on $\mathbb{R}_{+}$. The function $(x,a)\mapsto F_{x}(a)$ is $\bar{\mu}^{I}_{0}$-a.e. continuous while the function $\bar{\lambda}$ is $(\bar{\mu}_{X}\otimes Leb)$-a.e. continuous.
Regarding our next convergence results in probability, there is no loss of generality in considering the following assumption.
Assumption 3.3.
There exists $\omega^{*}>0$ that is an upper-bound of $\bar{\omega}^{N}$ uniformly on $\mathbf{\mathbb{X}}^{2}$ and on $N$.
@@
Note that the last identity defines $\mathbb{E}_{0}^{N}\big[\upsilon^{N}(i,j)\,\Big|\,(i,j)\in\mathcal{E}^{N}\big]$ as the ratio of $\mathbb{E}\big[\upsilon^{N}(i,j);(i,j)\in\mathcal{E}^{N}\,\big|\,\mathcal{F}_{0}^{N}\big]$ divided by $\mathbb{P}\big[(i,j)\in\mathcal{E}^{N}\,\big|\,\mathcal{F}_{0}^{N}\big]=\kappa^{N}(X_{i}^{N},X_{j}^{N})$.
For brevity, we use the notations $\mathcal{D}_{1}:=\mathcal{D}({\mathbb{R}}_{+},\mathcal{M}(\mathbf{\mathbb{X}}))$ and $\mathcal{D}_{2}:=\mathcal{D}({\mathbb{R}}_{+},\mathcal{M}(\mathbf{\mathbb{X}}\times\mathbb{R}_{+}))$.
@@
[…]
with $\bar{\lambda}$, $\bar{\omega}$ being given respectively in Assumptions 2.6 and 3.2. The uniqueness of the above system is stated among the potential candidates $(\bar{\mu}^{S}_{t},\overline{\mathfrak{F}}(t))_{t\geq 0}$ in which $\bar{\mu}^{S}_{\cdot}$ is a $\mathcal{M}(\mathbf{\mathbb{X}})$-valued càd-làg process such that $\mu_{t}(\mathbf{\mathbb{X}})\leq 1$ for any $t\geq 0$, while $\overline{\mathfrak{F}}(\cdot)$ is a càd-làg process whose values are bounded measurable functions from $\mathbf{\mathbb{X}}$ to $\mathbb{R}_{+}$, with local in time upper-bounds.
Given the pair $(\bar{\mu}^{S}_{\cdot},\overline{\mathfrak{F}}(\cdot))$ and the initial conditions $\bar{\mu}^{I}_{0}$ and $\bar{\mu}^{R}_{0}$, the explicit formulas for $\langle\bar{\mu}^{I}_{t},\psi\rangle$ and $\langle\bar{\mu}^{R}_{t},\phi\rangle$ with the test functions $\psi\in C_{b}(\mathbf{\mathbb{X}}\times{\mathbb{R}}_{+})$ and $\varphi\in C_{b}({\mathbb{R}}_{+})$ are as follows:
[…]
@@
$\begin{split}\langle\bar{\mu}^{R}_{t},\varphi\rangle&=\langle\bar{\mu}^{R}_{0},\varphi\rangle+\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\varphi(x)\bigg(1-\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bigg)\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)\\ &\quad+\int_{0}^{t}\hskip-5.69046pt\int_{\mathbf{\mathbb{X}}}\varphi(x)F_{x}(t-s)\overline{\mathfrak{F}}(s,x)\bar{\mu}^{S}_{s}(\mathrm{d}x)\mathrm{d}s\,.\end{split}$ (3.8)
Note that the unique solution to equation (3.5) is given by
$\bar{\mu}^{S}_{t}(\mathrm{d}x)=\,\exp\bigg(-\int_{0}^{t}\overline{\mathfrak{F}}(s,x)\mathrm{d}s\bigg)\bar{\mu}^{S}_{0}(\mathrm{d}x)\,$, (3.9)
and that (3.7) evaluated for any $x\in\mathbf{\mathbb{X}}$ with the function $\psi_{(x)}:(x^{\prime},a^{\prime})\mapsto\omega(x,x^{\prime})\frac{\bar{\lambda}(x^{\prime},a^{\prime})}{F_{x^{\prime}}^{c}(a^{\prime})}$ results in the following expression of $\overline{\mathfrak{F}}$ adapted from (3.6):
$\overline{\mathfrak{F}}(t,x)=\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\frac{\bar{\lambda}(x^{\prime},a^{\prime})}{F^{c}_{x^{\prime}}(a^{\prime})}\bar{\mu}_{t}^{I}(\mathrm{d}x^{\prime},da^{\prime})\,.$ (3.10)
A reformulation of this system of integro-delay equations (3.5)-(3.8) is proposed in Section 4.3 as a system of partial differential equations.
We also remark that in the special case of $\lambda_{i}^{N}(a)$ discussed in Remark 2.8, given the expression of $\bar{\lambda}(x,a)$ in (2.3), we obtain the following:
$\overline{\mathfrak{F}}(t,x)=\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\widetilde{\lambda}(x^{\prime},a^{\prime})\bar{\mu}_{t}^{I}(\mathrm{d}x^{\prime},da^{\prime})\,.$
Discussions of the main result
@@
Our conditions on the limiting kernel of interaction $\bar{\omega}$ encompasses most, if not all, of the (known) pairwise kernel interactions found in applications. Contact matrices over a discrete space are allowed, as, e.g., in [11], as well as spatial interactions as considered in [34] with characteristics that are distributed on a compact subset of ${\mathbb{R}}^{d}$, or even a combination of both settings. A possible extension of our result would be to consider individual types that evolve in time, typically when the individual locations move in space. Another natural objective would be to establish the FLLN in instances where the limiting kernel $\bar{\omega}$ incorporates an additional dependency on the overall distribution $\bar{\mu}_{X}$ as in the general setting of [34].
@@
For clarity like in Remark 2.11, let us assume for this presentation that $\bar{\kappa}^{N}=\bar{\kappa}$ for any $N$ and $\bar{\kappa}(x,x^{\prime})=x\cdot x^{\prime}$ while the $(X_{i}^{N})$ are uniformly distributed in $[0,1]$. Let us assume in addition that the contact rate is deterministic once the pairs of contacts are decided and that the contact rate function $\gamma^{N}$ does not depend on the types of the individuals $i$ and $j$ in contact while $\bar{\omega}$ is non-zero. In this case, $\Upsilon^{N}\equiv 0$.
@@
$\overline{\mathfrak{F}}(t)=\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}x^{\prime}\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime})}{F^{c}_{x^{\prime}}(a^{\prime})}\bar{\mu}_{t}^{I}(\mathrm{d}x^{\prime},da^{\prime})\,.$
As considered in [35], such a setting for the dynamics on a graphon can be generalized by assuming a scaling factor $\varepsilon^{N}>0$ depending on $N$, for instance $\varepsilon^{N}=N^{-\alpha}$ with $\alpha\in(0,1)$, in that $\bar{\kappa}^{N}=\varepsilon^{N}\cdot\bar{\kappa}$. Provided we again assume that the contact rate is fixed at some value $\bar{\gamma}^{N}>0$, we can follow the same reasoning except that $\bar{\gamma}^{N}$ is then equivalent to $\bar{\gamma}/(N\varepsilon^{N})$ instead of $\bar{\gamma}/N$ and that $N\varepsilon^{N}\to\infty$ is required (and sufficient) for Assumption 3.4 to hold.
Note also that $N\varepsilon^{N}\to\infty$ typically means that the node degrees are asked to go to infinity, without any condition on the relation to the number $N$ of nodes. We refer to [18] for more details on the relation to the denseness level of the graph (in terms notably of the number of edges or the node degrees). The additional condition proposed in [35] corresponding to the convergence to infinity of $N\varepsilon^{N}/\log(N)$ appears related to their technique of proof rather than to the convergence of the epidemic process itself. The case where $\varepsilon^{N}=N^{-1}$ leads to limiting equations of a different nature, as hinted by the exploratory simulations in [18] and clarified in the results of [10] for specific cases of stochastic block models.